OpenStudy (anonymous):

The first two terms of a trinomial are given. What number must be added to make the expression a binomial square? a) x^2 - 10x+____ b) x^2 + 11x+____

OpenStudy (anonymous):

For the first one, notice a =1 b = -10 To complete the square we need (b/2)^2 Using the order of operations b = -10 b/2 = -5 (b/2)^2 = 25 <------------ answer Can you try the second one?

OpenStudy (mertsj):

1/2 of 10 = 5 and 5^2 = 25 so add 5 to the first one

OpenStudy (mertsj):

1/2 of 11 is 11/2 and (11/2)^2 = 121/4. So add 121/4 to the second one.

OpenStudy (anonymous):

so do I divide the second number by 2? 10x and 11x

OpenStudy (anonymous):

yes, and then take that answer and times it by itself

OpenStudy (anonymous):

thx guys

OpenStudy (mertsj):

Not the x, just the coefficient

OpenStudy (anonymous):

yeah ok